A farmer sell a mixture of wild bird feed in 50 pound bags. If 15% of the mixture is sunflower seeds, how many pounds of sun flower seeds will he need to make 100 bags?
Answer by RAY100(1637)

# of sunflowers = .15 * 50 * 100 = 750 #
.
check ,,,,total wt is 100 bags * 50 # = 5000 #
.
% = 750 / 5000 = .15,,,,ok